Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 455     Entry time: 22 Nov 2023, 23:18
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 16.9 keV, up to 30 pulses and 1.1 MHz for alignment, 1 pulse for data collection
  • 700-800 uJ
  • stable delivery

Optical laser delivery

  • None

Achievements/Observations

  • More data collection with Jungfrau at the back of the CSS
  • Crystal diffraction data from salicylic acid, ibuprofen and MOF collected with detector close up

Issues

  • Clear aberations from our CRLs, so we operate with as few as possible
  • Twice during data collection we lost connection to JF4M module 01 requiering reset of receivers.
ELOG V3.1.4-7c3fd00